### v4

# 4.1. Degenerate and non-degenerate components Let us use the notation in (3.3) for the Littlewood-Paley partition of unity. Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing

4.1. Degenerate and non-degenerate components
Let us use the notation in (3.3) for the Littlewood-Paley partition of unity. Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then
In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing
Using the Littlewood-Paley decomposition for the Fourier integral operator we have
where the inequality with is justified because the symbol is supported on Now, we decompose with
where, as above,
and
where Here, is the “curvature” associated to the real-valued phase function Since sometimes we also write
for simplicity.
Note then that is given by
4.1.1. Symbol classes in projective co-ordinates
The aim of this subsection is to analyse the behaviour of symbols in the frequency variable with respect to the projective co-ordinates defined via
Although and are equal in value, the radial derivative keeps fixed, and the vertical derivative keeps fixed. In the next lemma we analyse the derivatives of a symbol in terms of the radial and of the angular variables. We note that and we write when and also for a row vector.
Lemma 4.1.
Let us assume that the domain of contains an open neighborhood of the region where is fixed. Assuming smooth in we have that
for some family of coefficients . Also,
on
Proof.
Let us prove (4.7). When
the chain rule gets
When observe that
Mathematical induction on
implies that can be written in terms of some coefficients as
For the proof of (4.10)
observe that assuming this identity valid for we have
With we apply (4.9) to deduce that
for some family of coefficients The proof of (4.7) is complete. For the proof of (4.8) note that for any
which verifies (4.8) when Assuming (4.8) with fixed, namely, that note that for any
from where the mathematical induction proves for all
For let be the class of symbols that consists of all functions which are Borel measurable in supported in the cone bundle
and satisfying the Hörmander type estimates
Lemma 4.2.
Let and let Let Consider the symbol
supported in the region
Then, we have the following symbol estimates
for in the support of
Proof.
Using that
satisfies the estimates in (4.12) and the identity (4.7) we have that
Observing that, on the region we have proved that
By the Leibniz rule we have that
In view of (4.8),
we can estimate
Since the sequence belongs uniformly in to we have that
Now, let us analyse the behaviour of the cutt-off at the “curvature” term Observe that for any
Then, we have the equalities
There is such that
Consequently,
Thus, the mathematical induction gives the estimate
We have proved the estimates
and
Note that when we recover the information that has order Indeed, let Since note that the order of with respect to is If we replace the analysis above with the symbol instead of then takes the place of and we get the estimate
Therefore we have proved that
The estimates for the derivatives of in and in above, the inequality in (4.20) and the Leibniz rule (4.16) conclude the proof of (4.14).
4.4. Control of the error operator
In this subsection we prove that the error operator is bounded on The idea behind the proof of this fact can be understood from the calculus of Fourier integral operators. The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [46, Page 19, Eq. (16)]. So, intuitively, if would have had the form in (1.1), namely, one could compute an asymptotic expansion for the symbol of as follows
and satisfying the asymptotic expansion, see e.g. [27, Page 17],
and the symbols satisfying the estimates (under suitable conditions on the phase )
for any
Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [46, Page 17] to deduce the boundedness of even under the presence of the oscillating term
First, let us calculate the kernel of the error operator Note that
Consequently,
On the other hand, the kernel of is given by
In order to prove the boundedness of it suffices to prove the following estimate on each dyadic component
However, the integral inside of this -norm, can be written as follows
So, let us simplify the notation above by defining
and
Observe that in view of the Euler homogeneity identity
we can re-write the phase of the integral as follows
and then
The analysis above shows that in order to prove the boundedness of it suffices to prove the following estimate
To do so, we will estimate the integral on the set of points that contribute more information to the integrand. This set of points is determined by the region:
By following the argument in [46, Page 19] one can make the approximation
where the error term has derivatives satisfying the estimate
Denoting we have that
Here, the main integral to study is Indeed, arguing as in [46, Page 17] one can demonstrate that is a smoothing symbol. Indeed, one can estimate this integral over the region by inserting the cut-off in the integrand of Since on and in view of the identity
one has that
Repeated integration by parts in the -variable provides the factor but the differentiation of the term gives the factor which shows that this portion of the integral is for any
Note that
where
and
To evaluate we compute the integral with respect to using the distributional identity
as follows
Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off
In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
Hence
So, the only stationary points occurs when and the Hessian at this stationary point is given by
Consequently, on the support of we have the Taylor series expansion
Similarly, in view of the Taylor expansion
we can write
In view of the stationary phase principle
Consequently,
Because captures the first term of the Taylor expansion of the symbol around we have proved that is essentially a symbol of order Moreover, the factor in shows that
In consequence, the difference
has order where Now, what we want to prove is that satisfies the inequality order Since
since is a smoothing symbol the order corresponds to the order of the symbol However, in view of (4.56), one has the estimates
Since also contains the term the amplitude satisfies the estimates
The previous inequality can be written as
where
In this way, the previous estimate says that, omitting the errors is essentially a symbol of order
Now, we have prepared the way in order to use Proposition 6.1 of [46] for deducing the estimate (4.55). Indeed, our estimate in (4.55) is the analog of (19) in [46, Page 19] which is a consequence of [46, Proposition 6.1]. For completeness, we present such a proposition as follows.
Lemma 4.13.
If is a symbol satisfying estimates of the type
when then, we have the kernel estimate
So, note that for the validity of (4.55)
we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
Proof of the main Theorem 3.1.
With the notation in Subsection 4.1 the operator has been decomposed into its degenerate and non-degenerate components, and respectively. In Subsection 4.2 in Proposition 4.4 we have proved that the degenerate operator is bounded on As for the non-degenerate operator we have constructed in Subsubsection 4.3.1 a Fourier integral operator which is an average operator, of the same order that and a pseudo-differential operator of order zero. We have proved, that the factorisation approach introduced in Tao [46] also operates in our case, namely, that is bounded on (see Proposition 4.9) is of weak (1,1) type in view of the standard Calderón-Zygmund theory, and that the error operator is bounded on see Subsection 4.4. Since then is of weak (1,1) type. The proof of Theorem 3.1 is complete.

### v5

@@
-4.1.
-Degenerate and non-degenerate components Let us use the notation in (3.3) for the Littlewood-Paley partition of unity.
-Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing
+The weak (1,1) boundedness of Fourier integral Operators with complex phases
+Abstract.
+Let be a Fourier integral operator of order associated with a canonical relation locally parametrised by a real-phase function.
+A fundamental result due to Seeger, Sogge, and Stein proved in the 90’s, gives the boundedness of from the Hardy space into Additionally, it was shown by T.
+Tao the weak (1,1) type of .
+In this work, we establish the weak (1,1) boundedness of a Fourier integral operator of order when it has associated a canonical relation parametrised by a complex phase function.
+This result in the complex-valued setting, cannot be derived from its counterpart in the real-valued case.
+Key words and phrases:
+Fourier Integral Operator, Complex Phase, Weak (1,1) continuity2020 Mathematics Subject Classification:
+35S30, 42B20; Secondary 42B37, 42B35Contents
+1. Overview
+Motivated by the developments about complex canonical relations, see e.g.
+Laptev, Safarov and Vassiliev [30], Melin and Sjöstrand [31], and Martini and Müller [29], in this paper we establish the weak (1,1) boundedness of Fourier integral operators with complex phases and of order Our approach combines two methods.
+First, we adapt to our setting the factorisation procedure developed by Tao [47] in the proof of the weak (1,1) inequality of these operators in the setting of real-valued phase functions, and the other one came from the global parametrisation of the canonical relation associated to the operator with an appropriate complex phase as observed by the second author in [40], when extending to Fourier integral operators with complex phases the - boundedness result established by Seeger, Sogge and Stein in [45].
+To present our main Theorem 1.7 we are going to introduce some preliminaries.
+We consider with and we denote by
+the Fourier transform of an integrable function
+1.1. Outline
+When working with the real methods of the harmonic analysis one can find three substantial categories related to the boundedness properties of operators: maximal averages, singular integrals, and oscillatory integrals.
+In the latter category, oscillatory integrals provide a necessary tool for exploiting the geometric properties related to the notions of curvature and of orthogonality.
+As it was pointed out by Stein in [44], this family of objects is not easily classified and appears in multiple forms: Bochner-Riesz operators, variants of pseudo-differential operators and of Fourier multipliers, and in the main object treated here, Fourier integral operators.
+Fourier integral operators appear as solutions to hyperbolic differential equations and then their boundedness properties have a direct relation with the validity of a priori estimates for these equations.
+As it was shown e.g.
+in [40, 45, 47], the geometry and the singularities of the canonical relation of an operator are reflected in the boundedness properties of operators in - spaces.
+Due to the compatibility of the space with the Littlewood-Paley theory, see Fefferman and Stein [17], and with the complex interpolation of operators, a main step in the proof of the -boundedness properties of Fourier integral operators (and for other operators arising in harmonic analysis), is the proof of a suitable criterion for its --boundedness which is usually interpolated with its corresponding -theory.
+However, as in the case of pseudo-differential operators, once proved the --continuity criterion one expects that the corresponding class of --bounded Fourier integral operators must satisfy the - boundedness property.
+This last property is called the weak (1,1) inequality/boundedness for these operators.
+The purpose of this paper is to prove the weak (1,1) boundedness for the class of Fourier integral operators associated to symbols of order and of complex phases (our main result is Theorem 1.7 below).
+This order is expected from the weak (1,1) inequality for Fourier integral operators with real-valued phase functions proved by Tao in [47], where it was applied a ‘second decomposition’ on the symbol of the operator adapted to a suitable family of ellipsoids with a smooth variation of eccentricity and orientation that avoid Kakeya-type covering lemmas, see Theorem 1.2.
+This approach was combined with a factorisation strategy.
+The order also provides bounded Fourier integral operators with real-valued phase functions from the Hardy space into , see Segger, Sogge, and Stein [45], as well as in the case of complex phase functions as proved by the second author in [40], see Theorems 1.3 and 1.6, respectively.
+The extension of the theory of Fourier integral operators due to Hörmander and Duistermaat to the case of complex-valued phase functions was systematically developed by Melin and Sjöstrand in [31] motivated by the problem of the construction of parametrices, or fundamental solutions for operators of principal type with non-real principal symbols, see for instance [32].
+The algebraic properties for the theory of Fourier integral operators with complex phases are well developed but their regularity has not been much studied, with the exception, e.g.
+of [40].
+In some sense, the use of complex phases provides a more natural approach to Fourier integral operators.
+Indeed, when working with the complex phase approach, the geometric obstructions of the global theory with real phase functions can be avoided and it is a remarkable fact that every Fourier integral operator with a real phase can be globally parametrized by a single complex phase, see e.g.
+[40].
+We also refer to Laptev, Safarov, and Vassiliev [30] for this construction.
+1.2. Fourier integral operators with real-phase functions
+In order to present our main result, we record the definition of Fourier integral operators based on the real-phase function approach.
+So, we present the following simplified version which is suitable for our purposes.
+Definition 1.1.
+A Fourier integral operator of order and with real phase function is a continuous linear operator of the form
+where is the Fourier transform of The symbol is of order that is satisfies the symbol inequalities
+The symbol has compact support in We assume that in an open neighbourhood of (with being a compact set) the real-valued phase function is homogeneous of degree in smooth in and satisfies the non-degeneracy condition
+The following result provides the weak(1,1) inequality for a Fourier integral operator of order or equivalently its boundedness from into (the weak -space)
+Theorem 1.2 (Tao [47], 2004).
+Let be a Fourier integral operator of order with a real phase function. Then is of weak type.
+The corresponding -theory for these operators as developed in [45], can be summarised as follows.
+Theorem 1.3 (Seeger, Sogge and Stein [45], 1991).
+Let Let be a Fourier integral operator of order with a real phase function. Then is bounded from to When then is bounded from to
+The case in Theorem 1.3 was proved by Duistermaat and Hörmander [15]111See also Èskin [16].
+using the -method and extended independently to the setting of Fourier integral operators with complex phases by Hörmander in [25] and by Melin and Sjöstrand in [32].
+One of the fundamental differences when working on the boundedness of operators in the setting of complex-valued phase functions, compared with the real-valued setting is that one has to deal with the boundedness properties of pseudo-differential operators in the class One reason for this is the analysis of the oscillatory term see Melin and Sjöstrand [32, Page 394].
+1.3. Fourier integral operators with complex-phase functions
+Let and be smooth manifolds of dimension and let us denote by the almost analytic continuation of For this notion see Hörmander [22] and Nirenberg [35], see also Subsection 2.2. Let
+be a smooth positive homogeneous canonical relation parametrised by a phase function of positive type, namely, it satisfies see Subsections 2.2 and 2.4 for details.
+The class of Fourier integral operators is determined modulo by those integral operators that locally have an integral kernel of the form
+where is a symbol of order By the equivalence-of-phase-function theorem, we can always microlocally assume that and that the symbol satisfies the Kohn-Nirenberg type estimates
+locally uniformly in According to Lemma 2.1 of [32], one can always re-write the kernel in (1.4) as
+where the function is of positive type.
+Note that the amplitude in (1.6) is probably different from the one in (1.4), however, we keep the same notation for simplicity.
+To be able to use the Fourier inversion formula we will assume that is independent of and compactly supported in So, we will work with symbols instead of amplitudes, in order to recover the simplified form
+To analyse the boundedness of a Fourier integral operator associated to the phase we require the following local graph condition.
+Assumption 1.4.
+There exists such that
+Locally, this means that
+on the support of the symbol for all
+Remark 1.5.
+According to Lemma 1.8.1 of [40], the local graph condition (1.8) in Assumption 1.4 is equivalent to the same assumption for almost every Moreover, the determinant in (1.9) is a polynomial function in which means that if (1.8) holds for some then it holds for all but finitely many To summarise this discussion, Assumption 1.4 is equivalent to the existence of the same condition for some, and then for almost every
+Before presenting our main result we record the -theory of Fourier integral operators with complex phases as developed by the second author in [40, Theorems 1.7.1 and 1.7.2] including the endpoint case
+Theorem 1.6.
+Let be a Fourier integral operator of order associated with a canonical relation parametrised by a complex phase of positive type and satisfying the local graph condition (1.8).
+If then is bounded from to and for is bounded from to
+1.4. Main result
+The following is the main result of this paper. We will assume that the complex phase function is of positive type, in such a way that when
+Theorem 1.7.
+Let be a Fourier integral operator of order associated with a canonical relation parametrised by a complex phase of positive type in such a way that when and satisfying the local graph condition (1.8).
+Then is of weak type.
+Remark 1.8.
+Below we briefly discuss some aspects of the main differences between the regularity properties of Fourier integral operators in the setting of complex-valued phase functions in comparison with the setting of real-valued phases.
+Remark 1.9.
+There is a simple argument to show that our main Theorem 1.7 does not follow from its real counterpart in Theorem 1.2. Indeed, any Fourier integral operator in the form
+with real and with real phase function
+has the symbol
+in the class if see Melin and Sjöstrand [32, Page 394]. Note that is bounded on with if, for example,
+However when this order is not good enough as the one in Theorem 1.2, or compared with the order for the --boundedness proved by Seeger, Sogge and Stein in [45] for real phase functions or even as proved in [40] for complex phase functions.
+In order to improve this order in Theorem 1.7 we need to avoid the reduction to a real phase function, making some strategic modifications to the argument given in Tao [47], in particular taking into account that we have to deal with dyadic components of the symbol
+Remark 1.10.
+We remark that in Tao [47] it has been introduced a factorisation approach for the proof of the weak (1,1) inequality of Fourier integral operators, see Section 4 for details.
+When has a real-valued phase function, according to such an approach a portion of the operator, the non-degenerate part is factorised as modulo an error operator Here, is a pseudo-differential operator of order zero and its symbol is essentially when where is the symbol of Note that the weak (1,1) boundedness of is a consequence of the Calderón-Zygmund theory.
+However, compared with the case of Fourier integral operators with complex phases, one has to redefine the operator because in this setting for Nevertheless, according to the -theory for the Hörmander classes see Álvarez and Hounie [1], a pseudo-differential operator with a symbol in these classes is of weak (1,1) type when In order to avoid the loss of regularity caused by the order , we will keep the symbol of the operator in the pseudo-differential part and we will absorb the oscillating term into the average operator Essentially, this term will modify the argument in [47], where the -boundedness of the averaging operator depends on the uniformly -boundedness of the Littlewood-Paley operators, but even, perturbating this part with the oscillating symbol we are able to preserve the -boundedness of In order to follow this strategy, we will make a fundamental reduction of our Theorem 1.7 to the case of phase functions of strictly positive type in Section 3, see Theorem 3.1.
+We present the proof of Theorem 1.7 in Section 3.
+The authors thank Terence Tao for suggesting the factorisation approach used in this manuscript.
+In the next Section 2, we present the preliminaries and the notation used in this work.
+We will continue the discussion about the regularity properties of Fourier integral operator in Section 3, its main result is Theorem 3.1 and its proof is presented in Section 4.
+We finish this manuscript with Section 5 where we present an application of our main theorem and also we provide a bibliographic discussion.
+2. Preliminaries
+Let and be paracompact smooth real manifolds.
+We assume that and are manifolds of dimension Here, we can take or When is defined in the usual form, namely, where denotes the Lebesgue measure on The measure of a Borel set is denoted by ans its characteristic function is denoted by or by We also consider the weak -space defined by the seminorm
+If is a compact manifold without boundary we define the spaces and by using partitions of unity (subordinated to local coordinate systems).
+We say that an operator is of weak (1,1) type if there exists a constant such that for all the inequality holds.
+A prototype of operators of weak (1,1) type is the family of Calderón-Zygmund singular integral operators, see Calderón and Zygmund [8].
+Below we precise the notation and the fundamental notions related to the theory of Fourier integral operators of importance for this work.
+We start with the properies of conic Lagrangian manifolds in the next subsection.
+2.1. Basics on symplectic geometry
+Now, let us follow [40, Chapter I].
+We recall that a -form is called symplectic on if and for all the bilinear form is antisymmetric and non-degenerate on .
+The canonical symplectic form on is defined as follows.
+Let be the canonical projection.
+For any let us consider the linear mappings
+The composition that is defines a 1-form on With the notation above, the canonical symplectic form on is defined by
+Since is an exact form, it follows that and then that is symplectic. If it follows that Now we record the family of submanifolds that are necessary when one defines the canonical relations.
+Let be of dimension A submanifold of dimension is called Lagrangian if
+We say that is conic if implies that for all
+Let be a smooth submanifold of of dimension Its conormal bundle in is defined by
+The following facts characterise the conic Lagrangian submanifolds of
+Let be a closed sub-manifold of dimension Then is a conic Lagrangian manifold if and only if the 1-form in (2.1) vanishes on
+Let be a submanifold of dimension Then its conormal bundle is a conic Lagrangian manifold.
+The Lagrangian manifolds have the following property.
+Let be a conic Lagrangian manifold and let
+(2.3) have constant rank equal to for all Then, each has a conic neighborhood such that
+1.
+is a smooth manifold of dimension
+2.
+is an open subset of
+1.
+The Lagrangian manifolds have a local representation defined in terms of phase functions that can be defined as follows.
+For this, let us consider a local trivialisation where we can assume that is an open subset of
+Definition 2.1 (Real-valued phase functions).
+Let be a cone in A smooth function is a real phase function if, it is homogeneous of degree one in and has no critical points as a function of that is
+Additionally, we say that is a non-degenerate phase function in if for any such that one has that
+is a system of linearly independent vectors on .
+The following facts describe locally a Lagrangian manifold in terms of a phase function.
+Let be a cone in and let be a non-degenerate phase function in Then, there exists an open cone containing such that the set
+(2.6) is a smooth conic sub-manifold of of dimension The mapping
+(2.7) is an immersion. Let us denote
+Let be a sub-manifold of dimension Then, is a conical Lagrangian manifold if and only if every has a conic neighborhood such that for some non-degenerate phase function
+Remark 2.2.
+The cone condition on corresponds to the homogeneity of the phase function.
+Remark 2.3.
+Although we have presented the previous Definition 2.1 of non-degenerate real phase function in the case of a real function of the same can be defined if one considers functions of Indeed, a real-valued phase function homogeneous of order 1 at that satisfies the following two conditions
+is called non-degenerate.
+2.2. Almost analytic continuation of real manifolds
+To define the almost analytic continuation of a manifold we require some preliminaries.
+We record that a function defined on an open subset is called almost analytic in if satisfies the Cauchy-Riemann equations on that is, if and all its derivatives vanish in Here, as usual,
+One defines an almost analytic extension of a manifold requiring that the corresponding coordinate functions are almost analytic in Here, we are going to present this notion carefully.
+Let and let be a non-negative and Lipschitz function.
+A function is called -flat in if for every compact set and for every integer one has the growth estimate for all This notion defines an equivalence relation on the space of mapping on
+If is a compact set, we will say that is flat on if it is -flat with One has the following property:
+let be -flat. Then all its derivatives are -flat and is flat on if and only if for all and all
+Let be an open subset of and let be a closed subset of A function is almost analytic on if the functions are flat on for all For an open set we will denote
+One can make the identification Note that every function defines an equivalence class of almost analytic functions on which consist of functions in which are almost analytic in modulo functions being flat on Any representative of this class is called an almost analytic continuation of in
+Let be an open subset of Let be a smooth submanifold of codimension of and let be a closed subset of Then, is called almost analytic on if for any point there exists an open set and -functions such that every is almost analytic on and such that in is defined by the zero-level sets with the differentials being linearly independent over
+Two almost analytic submanifolds and of are equivalent if they have the same dimension and the same intersection with namely, that
+and locally are flat functions on where and are the functions that define and respectively.
+A real manifold defines an equivalence class of almost analytic manifolds in A representative of this class is called an almost analytic continuation of in Now, let us consider:
+a real symplectic manifold of dimension and let be (modulo its equivalence class) its almost analytic continuation in
+Let be an almost analytic sub-manifold containing the real point and let be its real symplectic coordinates in a neighbourhood of
+Let be almost analytic continuations of the coordinates in in such a way that maps diffeomorphically on an open subset of
+Let be an almost analytic function such that in and such that is defined in a neighbourhood of by the equations
+Then, an almost analytic manifold satisfying this property, in some real symplectic coordinate system at every point is called a positive Lagrangian manifold.
+We conclude this discussion with the following result, see e.g.
+Theorem 1.2.1 in [40].
+Proposition 2.5.
+Let and be as before.
+If is another almost analytic continuation of coordinates in and is defined by the equation in a neighbourhood of then is locally equivalent to the manifold where is an almost analytic function and
+2.3. Fourier integral operators with real-valued phases
+We can assume that are open sets in . One defines the class of Fourier integral operators by the (microlocal) formula
+where the symbol is a smooth function locally in the class with This means that satisfies the symbol inequalities
+for in any compact subset of and while the real-valued phase function satisfies the following properties:
+1.
+for all ;
+2.
+3.
+is smooth (e.g. implies are linearly independent).
+Here is a Lagrangian manifold locally parametrised by the phase function
+Remark 2.6.
+The canonical relation associated with is the conic Lagrangian manifold in , defined by In view of the Hörmander equivalence-of-phase-functions theorem (see e.g.
+Theorem 1.1.3 in [40, Page 9]), the notion of Fourier integral operator becomes independent of the choice of a particular phase function associated to a Lagrangian manifold Because of the diffeomorphism we do not distinguish between and by saying also that is the canonical relation associated with
+2.4. Fourier integral operators with complex-valued phases
+Now we record the following result for canonical relations, see e.g. Theorem 1.2.1 in [40].
+Proposition 2.7.
+Let be a phase function of the positive type that is defined in a conical neighbourhood. Let be an almost analytic homogeneous continuation of in a canonical neighbourhood in Let
+Then, the image of the set under the mapping
+is a local positive Lagrangian manifold.
+Let and be smooth manifolds of dimension and let us denote by the almost analytic continuation of Let
+be a smooth positive homogeneous canonical relation. This means that is locally parametrised by a complex phase function satisfying the following properties:
+1.
+is homogeneous of order one in for all
+2.
+has no critical points on its domain, i.e.
+3.
+is smooth (e.g. implies that are independent over ),
+4.
+When the last condition is satisfied one says that the complex phase function is of positive type.
+Then, the class of Fourier integral operators is determined modulo by those integral operators that locally have an integral kernel of the form
+where is a symbol of order By the equivalence-of-phase-function theorem, we can always assume that and that the symbol satisfies the type -estimates
+locally uniformly in
+2.5. Asymptotics for oscillatory integrals
+Now, we recall two principles used in this work: the principle of non-stationary phase and the principle of stationary phase, respectively.
+Proposition 2.8 ([44], Page 342).
+Suppose that is smooth and supported in the unit ball, also let be real-valued so that for some multi-index with we have that
+throughout the support of Then,
+The constant is independent of and of and remains bounded as long as the norm of remains bounded.
+Proposition 2.9 ([49], Proposition 6.4).
+Let be a smooth function and assume that Assume also that the Hessian of at is an invertible matrix.
+Let be the signature of and let Let be supported in a small neighbourhood of Then for any the integral
+satisfies the asymptotic expansion
+where are differential operators of order with smooth coefficients depending on and bounds for finitely many derivatives of
+3. Reduction to phases of strictly positive type
+We say that a complex phase function is of strictly positive type, if when The aim of this section is to show that our main Theorem 1.7 follows from the following reduced version.
+Theorem 3.1.
+Let be a Fourier integral operator of order associated with a canonical relation parametrised by a complex phase of strictly positive type, that is, there exists such that for all one has the lower bound
+and additionally satisfying the local graph condition (1.8). Then is of weak type (1,1).
+The proof of Theorem 3.1 will be addressed in Section 4. Assuming this statement we give a proof of our main theorem.
+Proof of Theorem 1.7.
+Since the compactness of the closed annulus and of imply that there exists such that for all one has the lower bound
+In view of Theorem 3.1 we end the proof. ∎
+Notation
+We consider the dimension and the parameter We will write and the constant will be independent of and probably it is not the same in different parts of the manuscript.
+Always we will keep in mind that the constant can be chosen small enough by allowing to be arbitrarily small.
+Since we can define the projective system of co-ordinates via
+In the same way we will denote
+when
+We will use the standard notation for Littlewood-Paley decompositions. So, we fix a non-negative radial test function supported on the ball and for any we define the functions
+Even, when these functions are defined on we keep the same notation.
+We also will denote by the segment joining two arbitrary points
+4. Proof of Theorem 3.1
+The aim of this section is to prove the reduced version of our main Theorem 1.7, namely, the one in Theorem 3.1 for phase functions of strictly positive type.
+To do this we will combine the following two approaches:
+The first one, is a strategy traced back to the works of Melin and Sjöstrand [31, 32], where any Fourier integral operator can be written in the form
+(4.1) with the real parameter as in the local graph condition (1.8), the real phase function is given by
+(4.2) and the symbol
+belongs to the class This formula for the operator has shown to be effective when dealing with its --boundedness, see [40].
+The second one, introduced by Tao [47] where for the proof of the weak (1,1) inequality one decomposes the operator into its degenerate and non-degenerate components, and Indeed, this is done using the curvature notion, which measures the extent to which the phase function fails to be linear.
+The portion where the curvature is small on each dyadic component of the operator defines the degenerate component of , and its ‘complement’ defines the non-degenerate component Among other things, the following are the key points in the factorisation approach in [47]: (i) to prove that is bounded on ; (ii) to construct a pseudo-differential operator of order zero, and then of weak (1,1) type, according to the Calderón-Zygmund theory and an average operator which is bounded on and essentially has the same phase function that in such a way that the error operator is bounded on
+The aim of the next subsections is to show that the two previous approaches are compatible, allowing to factorise the operator with a complex phase satisfying the conditions in Theorem 3.1, into more manageable operators.
 4.1. Degenerate and non-degenerate components
@@
 where Here, is the “curvature” associated to the real-valued phase function Since sometimes we also write
-for simplicity.
-Note then that is given by
+for simplicity. Note then that is given by
 4.1.1. Symbol classes in projective co-ordinates
@@
 Proof.
-Let us prove (4.7). When
-the chain rule gets
+Let us prove (4.7). When the chain rule gets
 When observe that
-Mathematical induction on
-implies that can be written in terms of some coefficients as
-For the proof of (4.10)
-observe that assuming this identity valid for we have
+Mathematical induction on implies that can be written in terms of some coefficients as
+For the proof of (4.10) observe that assuming this identity valid for we have
 With we apply (4.9) to deduce that
@@
 which verifies (4.8) when Assuming (4.8) with fixed, namely, that note that for any
-from where the mathematical induction proves for all
+from where the mathematical induction proves (4.8) for all ∎
 For let be the class of symbols that consists of all functions which are Borel measurable in supported in the cone bundle
@@
 Proof.
-Using that
-satisfies the estimates in (4.12) and the identity (4.7) we have that
+Using that satisfies the estimates in (4.12) and the identity (4.7) we have that
 Observing that, on the region we have proved that
 By the Leibniz rule we have that
-In view of (4.8),
-we can estimate
+In view of (4.8), we can estimate
 Since the sequence belongs uniformly in to we have that
@@
 Therefore we have proved that
-The estimates for the derivatives of in and in above, the inequality in (4.20) and the Leibniz rule (4.16) conclude the proof of (4.14).
+The estimates for the derivatives of in and in above, the inequality in (4.20) and the Leibniz rule (4.16) conclude the proof of (4.14). ∎
+4.2. Boundedness of the degenerate component
+In this section we prove that the operator
+is bounded on Before continuing with the proof of the -boundedness of we discuss the second dyadic partition that will be employed in our further analysis.
+Remark 4.3 (About discrete vs continuous ‘second’ dyadic decompositions).
+In [45], a discrete ‘second dyadic decomposition’ was introduced where the variable is decomposed in a finite family of disks see Figure 1.
+At first glance, for the analysis of the degenerate component one could apply this discrete ‘second dyadic partition’ on the symbol:
+generating the new family of symbols in such a way that However, as it was explained in [47], this discrete decomposition provides the estimate
+for the -norm of the kernel associated with each dyadic part This estimate is not enough to deduce the boundeness of The reason is that one needs a more accurate decay of this -norm.
+Indeed, this is proved in Lemma 4.6 improving the term in (4.22), by the smaller one In order to make this improvement, we will use the degeneracy condition To have the control of lower terms in the Taylor expansion of the phase at especially if is in the direction where degenerates, the disks will be replaced by a suitable family of ellipsoids, that allows us to decompose the -variable in a family of tubular regions where we gain the decay
+As it was pointed out in [47], there is an apparent difficulty in the case when where the ellipsoid around depends on the spectrum of Then, the eccentricity and the orientation of this ellipsoid vary in function of However, by adapting to the setting of the complex phases, the approach in [47], we will define a continuous ‘second dyadic partition’ in using a family of functions with a family of corresponding supports determined by a family of ellipsoids with a smooth variation of eccentricity and orientation, centred at the points that avoid Kakeya-type covering lemmas.
+According to this partition the part of the support of each symbol on the hyperplane is illustrated in Figure 2.
+The following is the main result of this subsection.
+Proposition 4.4.
+Let be the symbol of the Fourier integral operator of order and let be its complex phase function. Then, the degenerate component in (4.21) extends to a bounded operator from into
+Remark 4.5.
+For the proof of Proposition 4.4 we require some preliminary results.
+We start with the following estimates for radial and angular derivatives of symbols in the class and with an estimate of the -norm of the kernel of This is our first step in the proof of (4.24).
+Lemma 4.6.
+Let and let Let Consider the symbol
+Then, we have the following symbol estimates
+for in the support of Moreover, any function supported in the region
+satisfying the symbol estimates in (4.27) defines a Fourier integral operator whose kernel satisfies the inequality
+uniformly in
+Proof.
+Now, let us consider a symbol satisfying the estimates in (4.27) whose support is contained in (4.28).
+For the proof of (4.29) we need a preliminary construction.
+Indeed, let us define a positive-definite replacement of Note that is the associated matrix to the quadratic form
+Let
+be the family of eigenvalues of In an appropriate bases of we can write
+Define
+where is the identity matrix of size In the same basis we can write
+Observe that
+On the other hand
+The matrix has determinant satisfying the estimates
+For any with the property that when
+define the function by
+Since on we necessarily have that and so is contained in a compact subset of Note that
+In consequence, if
+then
+and consequently contains a small disk namely, we have the inclusion
+By writing in the basis we have
+from where one can deduce that is supported in an ellipsoid centred at Observe that, since
+the support of contains the disk Indeed, on the support of which says that
+In view of (4.35) the inequality implies that
+In consequence,
+In view of the inclusions (4.34) and (4.36) for any we have
+Now, define the averaged function by
+We review the properties of the averaged function in the following lemma.
+Lemma 4.7.
+The function is positive, and the following estimates are valid:
+Proof.
+In a suitable basis we can write the quadratic form as follows
+Observe that
+By making Taylor expansion of the function at we have that
+The change of variables has the new volume element
+from where we have that
+A similar analysis can be done for the proof of the estimate On the other hand, let Note that when applying derivatives to the main term vanishes and the error term contribute in ∎
+In order to prove (4.29) we use the continuous partition of unity taking averages with respect to ellipsoids centered at as follows (and using Fubini’s theorem)
+The variable runs over a compact set when as discussed above. Since on the support of Figure 2 illustrates the points of the support of on the hyperplane
+Hence, for the proof of (4.29) is enough to show that
+Indeed, observe that (4.40) implies the estimate
+For the proof of (4.40) we write the frequency variable in terms of the co-variables and of the new volume element
+Then, (4.40) can be written as
+To simplify the previous estimate we will consider the following change of variables
+with a constant Note that
+In what follows, let us denote
+and
+Then, in view of (4.33) and (4.42) we have that
+Then, by multiplying in both sides of (4.41) by the factor we can re-write (4.41) in terms of as follows
+where
+So, to finish the proof we will show the estimate in (4.43). The advantage in proving (4.43) is that the symbol allows the use of the phase stationary theorem since it does satisfy the estimates
+Note that, when the chain rule gives
+where we have used that on the support of see (4.28). Then, using mathematical induction we can prove the estimate
+where the previous estimate depends on but it is independent of if is large enough and we take a finite number of derivatives in .
+This will be the case because we are going to apply the method of stationary phase just taking derivatives in with Now, using that the derivatives of are uniformly bounded, in order to prove that the derivatives in satisfy (4.44) we proceed as follows for
+since Note that we have the estimate
+Moreover, since is a polynomial of order when and we trivially have the bound
+In consequence, since we can estimate
+On the other hand, to complete the proof of (4.44) we just have to use the estimate
+Since we can estimate
+as claimed. Indeed, note that the derivatives with are bounded from above by when On the other hand, let us analyse the phase using its Taylor expansion as follows taking into account that
+Since the phase function
+has gradient
+we have that
+Then, with one has the estimate
+For our further analysis note that
+So, in order to apply the non-stationary phase principle222See Proposition 2.8. We are going to apply this proposition to the phase making use of the estimate (4.45). to the integral
+we re-write it as follows
+Consequently,
+where the eccentric disc is given by
+with Taking into account that one can estimate the volume of as follows, see [47, Page 8]
+So, we can estimate
+Taking small enough and large enough we have the estimate Therefore, we have that
+as desired. We have proved (4.43). In consequence, the proof of Lemma 4.6 is complete. ∎
+Proof of Proposition 4.4.
+4.3. Boundedness of the non-degenerate component
+Now, we are going to prove the weak (1,1) boundedness of the non-degenerate component in (4.5).
+To do this, we follow the approach in [47], by making the factorisation modulo an error operator bounded on Here, we are going to construct a pseudo-differential operator of order zero and an averaging operator which is bounded on
+4.3.1. Construction of the operators and
+Let us consider the projection in the first component and let be an open subset, such that is compact with Let be the characteristic function of The operator will be defined as follows
+where denotes the value of at and is a bump function to be chosen later.
+The integer quantities are such that they satisfy, according to the principle of stationary phase, the asymptotic expansion (see Proposition 2.9)
+when Since are smooth functions in the region of integration they are constant on each connected component of this region.
+Remark 4.8.
+When that is if is real-valued, the definition of purely involves the phase function and the curvature term the -boundedness of follows from the argument for real-valued phase functions in Tao [47, Pages 14-15].
+Lemma 4.9.
+extends to a bounded operator on provided that there exists such that for all the lower bound
+holds.
+Proof.
+First, assume that is smooth and compactly supported and that vanishes when Note that the support of is contained in the region By applying summation by parts, because of the properties of the support of boundary terms vanish and we can write
+In view of the triangle inequality, it suffices to prove the estimate
+where
+Let us denote by
+Then,
+Using the Fourier inversion formula, we have that
+Since is contained in the set where the difference is supported on the region Since we have that
+and consequently
+Note that the operator
+is uniformly bounded in on Indeed, its kernel given by
+belongs to uniformly in Since the phase function is of strictly positive type, that is, and note that
+Indeed, since when the compactness of the closed annulus and of imply that there exists such that for all one has the lower bound
+Therefore,
+Thus, we have proved that and Schur’s lemma gives the boundedness of on Note also, that if is an smooth change of coordinates on , the operator
+is also bounded on Moreover, since by following the previous argument, we have that
+Now, since for any is a diffeomorphism on taking large enough such that we have that
+Since is in and is compact, So we have the estimate
+as desired proving the -boundedness of The proof of Lemma 4.9 is complete. ∎
+Definition 4.10.
+Using that the mapping
+is a diffeomorphism on the support of the symbol we define the pseudo-differential operator associated to the symbol
+where, is a bounded function away from zero and smooth for all and for all .
+Remark 4.11.
+Since is an open neighbourhood of the set we have that on In consequence is smooth in and
+Now we prove the following fundamental property of the operator
+Lemma 4.12.
+The pseudo-differential operator defined by
+is of weak (1,1) type.
+Proof.
+Since and we have that is symbol of order zero. Then, the pseudo-differential operator is essentially a Calderón-Zygund operator and then it satisfies the weak (1,1) inequality, namely, is bounded. ∎
 4.4. Control of the error operator
 In this subsection we prove that the error operator is bounded on The idea behind the proof of this fact can be understood from the calculus of Fourier integral operators.
-The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [46, Page 19, Eq.
+The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [47, Page 19, Eq.
 (16)].
@@
 and the symbols satisfying the estimates (under suitable conditions on the phase )
-for any
-Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [46, Page 17] to deduce the boundedness of even under the presence of the oscillating term
+for any Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [47, Page 17] to deduce the boundedness of even under the presence of the oscillating term
 First, let us calculate the kernel of the error operator Note that
@@
 To do so, we will estimate the integral on the set of points that contribute more information to the integrand. This set of points is determined by the region:
-By following the argument in [46, Page 19] one can make the approximation
+By following the argument in [47, Page 19] one can make the approximation
 where the error term has derivatives satisfying the estimate
 Denoting we have that
-Here, the main integral to study is Indeed, arguing as in [46, Page 17] one can demonstrate that is a smoothing symbol.
+Here, the main integral to study is Indeed, arguing as in [47, Page 17] one can demonstrate that is a smoothing symbol.
 Indeed, one can estimate this integral over the region by inserting the cut-off in the integrand of Since on and in view of the identity
@@
 as follows
-Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off
-In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
+Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
 Hence
@@
 Now, we have prepared the way in order to use Proposition 6.1 of [47] for deducing the estimate (4.55).
-Indeed, our estimate in (4.55) is the analog of (19) in [46, Page 19] which is a consequence of [46, Proposition 6.1].
+Indeed, our estimate in (4.55) is the analog of (19) in [47, Page 19] which is a consequence of [47, Proposition 6.1].
 For completeness, we present such a proposition as follows.
@@
 when then, we have the kernel estimate
-So, note that for the validity of (4.55)
-we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
+So, note that for the validity of (4.55) we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
 Proof of the main Theorem 3.1.
@@
 The proof of Theorem 3.1 is complete.
+∎
+5. Final remarks
+5.1. Applications
+In order to illustrate the applications of Theorem 3.1 to a priori estimates for the Cauchy problem with complex characteristics we consider the case where the associated operator is a classical pseudo-differential operator of order one.
+Let us introduce some required notation.
+Let be a compact Riemmanian manifold of dimension and let us consider Denote by the positive Laplacian associated to the metric Denote the space defined by the norm
+We also consider the weak -space defined by the seminorm
+and by its local version, namely the family of functions such that for any compactly supported function
+Denote the partial derivative with respect to the time variable and let Let us consider the Cauchy problem for a first-order classical pseudo-differential operator
+We assume that its principal symbol has non-negative imaginary part, namely, satisfies
+The solution operator of (5.1) can be realised as a Fourier integral operator of order zero with a global complex phase function of positive type, see [40, Page 117] and also Laptev, Safarov and Vassiliev [30] for details.
+Applying (essentially) Theorem 3.1 to the operator
+of order we have that for any compactly supported initial datum , and for a fixed time the solution of the Cauchy problem (5.1) satisfies This analysis proves the following consequence of Theorem 3.1.
+Corollary 5.1.
+Let and let be compactly supported. Then, for a fixed , the solution of the Cauchy problem (5.1) satisfies
+Proof.
+We write namely, the solution of (5.1) in terms of the solution operator Since we have that
+We can use the factorisation
+and the fact that is a Fourier integral operator with complex phase of order 333See [40, Page 117] and also Laptev, Safarov and Vassiliev [30] for details.
+to prove that Indeed, if denotes a partition of unity subordinated to a finite open covering of the compact manifold then and We can assume that the covering is an atlas for the manifold and then each is diffeomorphic to an Euclidean open set Note that is a Fourier integral operator of order that can be microlocalised to a Fourier integral operator whose symbol is compactly supported in the spatial variable.
+Here, denotes the multiplication operator by the function Then, Theorem 3.1 imply that for any we have that
+which proves Corollary 5.1. ∎
+We also observe that a similar analysis can be adapted to investigate the mapping properties of the parametrix of other first-order complex partial differential operators, see e.g.
+[40, Page 120] for a problem that appears naturally in the study of the oblique derivative problem.
+This was observed e.g.
+by Melin and Sjöstrand in [31].
+For other applications of the theory of Fourier integral operators with complex phases to a priori estimates for the Cauchy problem with complex characteristics we refer to Trèves [48, Chapter XI] and to [40, Chapter 5].
+5.2. Conclusions and bibliographical discussion
+In this work we have analysed the weak (1,1) boundedness of Fourier integral operators with complex-valued phase functions.
+As in the setting of real-valued phase functions regarding the weak (1,1) boundedness due to Tao in [47], our estimate is local in the sense that we have considered the symbol to be compactly supported in the -variable, following the previous approaches, see Seeger, Sogge and Stein [45], Beals [3], and the monograph of the second author about the mapping properties of Fourier integral operators with complex phases [40].
+Among other things, the sharpness of the order for the weak (1,1) inequality of elliptic Fourier integral operators has been discussed by the authors in [9], and we also refer to [39] for the sharpness of Seeger-Sogge-Stein orders regarding the -boundeness of these operators.
+There has been in the last decades a wide interest in the analysis of the global continuity properties for Fourier integral operators.
+This problem can be traced back to the work of Asada and Fujiwara [2] and of Peral [36] and Miyachi [34].
+We refer to [41], [42] and [12, 13] for global properties on -spaces.
+As for global criteria for Fourier integral operators on Hörmander classes removing the calculus condition and allowing the ranges and we refer the reader to Dos Santos Ferreira and Staubach [14], Kenig and Staubach [28] and the works of Staubach and his collaborators [26, 27, 37, 38].
+We also refer to the seminal work of Álvarez and Hounie [1] for the boundedness of pseudo-differential operators without the condition which was the starting point for the aforementioned results in the setting of Fourier integral operators.
+For the global parametrization of Lagrangian distributions we refer the reader to Laptev, Safarov, and Vassiliev [30] and to [40].
+We also refer to the work [10] for the global definition of Fourier integral operators on compact Lie groups in terms of the group Fourier transform.
+The theory of Fourier integral operators with complex phases was developed by Melin and Sjöstrand in [31] motivated by the problem of the construction of parametrices for operators of principal type with non-real principal symbols, see for instance [32].
+The singularities of the Bergman kernel can be also approximated in terms of asymptotic expansions of certain Fourier integral operators with complex-valued functions, see Boutet de Monvel and Sjöstrand [7].
+For a systematic analysis of the Bergman kernel we refer to Fefferman [18, 19].
+There has been also wide activity concerning the smoothing estimates for Fourier integral operators after the proof of Bourgain and Demeter [6] of the decoupling conjecture.
+The smoothing effect for Fourier integral operators was conjectured by Sogge in [46].
+It will be difficult to review the literature about this problem here, but in what follows we share some works in this direction.
+Indeed, we refer the reader to Bourgain [5], Minicozzi and Sogge [33], Guth, Wang and Zhang [21], Beltran, Hickman and Sogge [4], to Gao, Liu, Miao, and Xi [20] and references therein for details.
+Conflict of interests statement - Data Availability Statements The authors state that there is no conflict of interest.
+Data sharing does not apply to this article as no datasets were generated or analysed during the current study.
+5.3. Acknowledgement
+The authors thank Terence Tao for suggesting the factorisation approach used in this manuscript.